OpenStudy (anonymous):

Assume that a company in a year from now will have a stream of profits £100 every year for the rest of time (i.e. forever). What is the present discounted value of an income stream of £100 when the interest rate is 3%?

OpenStudy (anonymous):

This company run by someone's grandma? Those are slim margins.

OpenStudy (anonymous):

Must be!

OpenStudy (anonymous):

Okay, to do this problem we have to work backwards.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

First of all, do you know any equations that apply here?

OpenStudy (anonymous):

Well, usually (when working with a single year) I'd do the following:

OpenStudy (anonymous):

100/(1+0.03)^1

OpenStudy (anonymous):

But in this case, that comes to £97.08, which I'm told is not the answer.. So I'm assuming I need to do this: 100/(1+0.03)^(infinity) ?

OpenStudy (anonymous):

What would you do for two years?

OpenStudy (anonymous):

100/(1+0.03)^1 + 100/(1+0.03)^2 ?

OpenStudy (anonymous):

Correct, assuming you get 100 each year.

OpenStudy (anonymous):

If so, I retract my earlier assumption of 100/(1+0.03)^(infinity) as it'd need to be a summation, right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So I need a way of working a summation up to the 100/(1+0.03)^(infinity) ?

OpenStudy (anonymous):

What we have is SUM ar^k = a SUM r^k = (a) / (1-r) when the SUM is infinite

OpenStudy (anonymous):

http://mathworld.wolfram.com/GeometricSeries.html

OpenStudy (anonymous):

Does this display properly for you: \[ \frac ab \]

OpenStudy (anonymous):

I see \[ \frac ab \]

OpenStudy (anonymous):

Okay unfortunately math on this site is broken

OpenStudy (anonymous):

Ah, okay! Well thank you for your help so far!

OpenStudy (anonymous):

Ah!

OpenStudy (anonymous):

Anyway, the point is: r + r^2 + r^3 + ... + r^infinity = r / (1 - r)

OpenStudy (anonymous):

Do you think you can do this problem knowing that?

OpenStudy (anonymous):

That's a big help. The only issue I'm having now is in thinking what should I treat as r. Could I work with r as 1.03 initially?

OpenStudy (anonymous):

Hm, definitely not right

OpenStudy (anonymous):

No, remember that r < 1

OpenStudy (anonymous):

100/(1+0.03)^k

OpenStudy (anonymous):

Another way to write this is 100*(1+0.03)^-k

OpenStudy (anonymous):

Which also is 100*[(1+0.03)^-1]^k

OpenStudy (anonymous):

So 100 + (1/3)^k ?

OpenStudy (anonymous):

No plus signs here.

OpenStudy (anonymous):

Identify a and r.

OpenStudy (anonymous):

Hint: a*[r]^k

OpenStudy (anonymous):

Really lost. Trying to get there but not I've just got 100's and 3's all over my page. I'll keep trying

OpenStudy (anonymous):

Wait, is a = 100 and r = 100/103 ?

OpenStudy (anonymous):

It is!! I think I'm there!! 100 * [(100/103)/(1- (100/103))]

OpenStudy (anonymous):

Thanks so much for your help - I really appreciate it!

OpenStudy (anonymous):

Did you get correct answer?

OpenStudy (anonymous):

I did!